a) A Numeracy 2 resit group consists of 22 male and 18 female students. What is the probability of their tutor picking at random that a female will be selected?
Answer by rothauserc(4718)

there are 22 + 18 = 40 students in the group
:
probability of selecting a female = 18/40 = 9/20
: